# Mitigating Forgetting in LLM Fine-Tuning via Low-Perplexity Token Learning

### v3

@@ B.1 STM performance on different models
Table 11: STM performance comparison between Gemma 2 IT 2B Instruct, Mistral 7B-Instruct and Llama3 8B Instruct fine-tuned on MATH target task. The percentage indicates the number of filtered tokens from the training data.
@@
Mistral 7B-Instruct
Ground Truth 16.7 47.1 0.0 35.0 16.2
@@
B.2 Optimal threshold selection for STM
@@
B.3 Alternative STM strategies of Llama 3 8B Instruct.
@@
B.4 Other fine-tuning techniques applied with STM on Llama 3 8B Instruct
@@
B.5 Convergence curve with Gemma 2 IT 2B
@@
Appendix C High perplexity tokens filtered by STM
We have noticed that the Self-output responses consist with low surprisal values as highly predictable token sequences. It is also worthwhile to note that the contents of high perplexity tokens between Ground Truth responses and Self-ouotput responses are different. In Math dataset, we simply categorize the tokens in model responses as Numbers, Symbols and Words. Numbers refers to the tokens which are numeric words. Symbols refers to tokens that imply a non-numeric symbols or math word in latex, such as +,,,/+,-,*,/+ , - , ∗ , /. Words refers to tokens consisting of alphabets only. As Figure 8, Figure 9, and Figure 10 shows, in terms of all the models we have tested, the proportion of number tokens with high perplexity drop drastically in Self-output responses, while the overall distribution of the three categories in both Ground Truth responses and Self-Output responses are almost the same. It implies that the number tokens in Ground Truth dataset are less predictable by models and harder to learn.
Figure 8: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth dataset and Self-Output dataset of MATH task with Llama3-8B-Instruct
Figure 9: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth dataset and Self-Output dataset of MATH task with Mistrak-7B
Figure 10: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth dataset and Self-Output dataset of MATH task with Gemma2-2B-it
Appendix D Reproducibility
D.1 Training data cost
As we know there is efficiency difference of the data preparation between Self-Output, Rephrase and Selective Token Masking/Ground Truth. To brief about the difference, Self-Output is the most resource-exhausted since it requires to generate $N$ samples (usually set 32), validate the correctness of samples (including parsing out the answer, and match with gold label or pass test cases), and then do forward passes on those samples to pick the lowest perplexity sentence as training data. On the other hand, STM only requires a single forward pass on Ground Truth data to calculate perplexity of each token to do the masking in training process. As Table 14 shows, for a simple MBPP task, on a A100 GPU server to prepare for 374 instances (train) + 90 instances (validation) for $N$=32 takes about more than 18 gpu hours to complete the process of Self-Output to generate trainable data. While STM only requires 4 minutes to calculate the whole token perplexity.
@@ D.1 Training data cost
Ground Truth 374 100% 0
STM Llama3-8B Insruct 374 100% 4.5 minutes
Self-Output Llama3-8B Instruct 213 57% 16.8 hours
Rephrase Llama3-8B Instruct 327 87.4% 36.8 minutes
D.2 Model training resource
We train Llama 3 8B Instruct, Mistral 7B-Instruct and Gemma 2 IT 2B using two NVIDIA A100 Tensor Core GPU with VRAM 40GB for each GPU and additional RAM of 96GB. In terms of gpu hour, Our training experiments takes at most 2 hours for MBPP task, 4 to 6 hours for MATH task. As for the evaluation experiments for other non-target tasks, evaluation in ARC-Challenge takes 3.5 hours, evaluation in GSM8k takes 4.5 hours, and evaluation in BIRD task takes 7.5 hours.
D.3 Evaluation prompt
@@
MBPP prompt user prompt Please refer the given test cases and generate a python function for my problem. Make sure the written code is wrapped in code block : ```python $<yourcode>$``` >>>much-greater-thanabsent>>>> > > Problem:{text}>>>much-greater-thanabsent>>>> > > Test Cases:{test_list}
@@
ARC Challenge prompt user prompt Answer the following multiple choice question. The last line of your response should be of the following format: ’Answer: \$$\$LETTER^{\prime}$ (without quotes) where LETTER is one of ABCD. Question: {question} Choices: {choices}
In GSM8k task, we add a Question-Answering task instruction and an answer format instruction before the ’question’ as the user role prompt. To evaluate the output, we directly parse the format of ’####$\#\#\#\#<number\,\,only\,\,answer>$’ from the generated results, and match it with the ’answer’ value from the dataset. The TextBox shows an example of the user prompt:
GSM8k prompt user prompt You are given a grade school math question. Please answer the question in the following format: Q: $<$question$>$ A: $<$Think step by step here$>$ #### $<$number only answer$>$ Format requirements : you must first output your reasoning before finalized with the " #### " format followed by the final numeric answer.
@@
Appendix E Other STM alternatives
@@
Base model 0% 28.1 51.1 58.5 74.7 13.9
Ground 0% 21.7 45.8 19.0 29.5 5.1
Truth
$STM$ 23.8% 30.3 49.5 55.4 75.3 13.3
@@
Appendix F Correctness experiment with other models
We have also tested the ablation test of correctness with Gemma2-2B-Instruct model as Figure 16 shows. It holds the same conclusion that correctness is relevant to in-domain performance only, while the low perplexity trait of Self-Output data is more relevant to out-of-domain performance than correctness. However in Gemma2-2B setting, the number of training instance with a least one positive and one negative responses at the same time is much fewer than ground truth data and Llama’s correctness data although it has more trainable tokens due to different tokenizers are applied. That is to say, to maintain the same training data amount between each correctness rate dataset, we have to choose a smaller training data size for Gemma2-2B. Therefore, the effectiveness of data correctness is not as obvious as Llama3-8B model does.
Table 16: The performances of models trained with different correctness rate of MBPP Self-Output dataset and trainable tokens number. Values of a cell block in gray represents the in-domain testing performance of the Gemma2-2B Instruct trained with in-domain task data.
@@ Appendix F Correctness experiment with other models
GT 86,632 .400 .193 .229 .716 .051
Self Output
100% 87,632 .500 .287 .742 .563 .134
49% 90,594 .489 .286 .732 .564 .134
0% 90,416 .497 .284 .742 .563 .135
base model - .511 .281 .747 .585 .139
Appendix G Regularization effects on non-target tasks
We apply a range for searching best combination of weight decay and dropout rate hyperparameters when training an MBPP task on Llama3-8B-Instruct model. The weight decay is set from 0.0 to 0.5 with interval 0.1, and dropout rate is set from 0.05 to 0.5 with an interval of 0.05 to 0.1. And we pick the top 3 best performing weight decay value with its dropout rate, and choose the best 2 performances among the checkpoints of 50 steps to 300 steps (about 5 to 6 epochs training) as our results. Our experiment results as Figure 11 show that, when applying dropout or weight decay on ground truth data, the model trained with target task like MBPP still overfits on training data and the testing performance degrades. On the other hand, the non-target performance of MATH as Figure 12 shows, we found that some of the results perform better than the original Llama in non-target task, but their target task performances are found worse than the base model’s performance. It implies that the widely adopted regularization techniques are hard to optimized for both target task and non-target task performance.
Figure 11: The target task performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MBPP testing dataset.
Figure 12: The non-target MATH task performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MATH testing dataset.
Appendix H Rank analysis on LoRA weights

### v6

@@ D.1 STM performance on different models
Table 13: STM performance comparison between Gemma 2 IT 2B, Mistral 7B Instruct and Llama3 8B Instruct fine-tuned on MATH target task. The percentage indicates the number of filtered tokens from the training data.
@@
Mistral 7B Instruct
Baseline Fine-tuning 16.7 47.1 0.0 35.0 16.2
@@
D.2 Optimal threshold selection for STM
@@
D.3 Alternative STM strategies of Llama 3 8B Instruct.
@@
D.4 Ablation and scaling on STM threshold for Llama 3 8B Instruct
We have done experiments of STM applying ablation and scaling of thresholds as Table 15 shows: high ppl threshold and and optimal threshold around 20%to 25% yields the best performance in terms of BWT and TI.
Table 15: Ablation and Scaling on STM threshold for Llama 3 8B Instruct model on MBPP target task.
Llama 3 8B Instruct BWT(%) TI(%)
$STM_{\tau=2.5,high}$ -0.16 3.2
$STM_{\tau=2.5,random}$ -40.41 -9.84
$STM_{\tau=2.5,low}$ -33.01 -14.73
Baseline Fine-Tuning -14.1 -17.8
$STM_{\tau=1000}$ (1.7%) -19.6 -11.8
$STM_{\tau=10}$ (11.0%) -13.1 5.1
$STM_{\tau=2.5}$ (22.2%) -0.3 6.4
D.5 Other fine-tuning techniques applied with STM on Llama 3 8B Instruct
@@
D.6 Convergence curve with Gemma 2 IT 2B
@@
Appendix E High perplexity tokens filtered by STM
We have noticed that the Self-output responses consist with low surprisal values as highly predictable token sequences. It is also worthwhile to note that the contents of high perplexity tokens between Ground Truth responses and Self-ouotput responses are different. In Math dataset, we simply categorize the tokens in model responses as Numbers, Symbols and Words. Numbers refers to the tokens which are numeric words. Symbols refers to tokens that imply a non-numeric symbols or math word in latex, such as +,,,/+,-,*,/. Words refers to tokens consisting of alphabets only. As Figure 8, Figure 9, and Figure 10 shows, in terms of all the models we have tested, the proportion of number tokens with high perplexity drop drastically in Self-output responses, while the overall distribution of the three categories in both Ground Truth responses and Self-Output responses are almost the same. It implies that the number tokens in Ground Truth dataset are less predictable by models and harder to learn. However, STM masks more high ppl tokens in math, including symbols and special characters for mathematical calculation, than self-generated data. But the performance of TI is still positive and better than baseline fine-tuning, showing that the model still improves without knowing those important key symbols but from the context without the masked words.
Figure 8: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth (GT)dataset and Self-Output (SO) dataset of MATH task with Llama 3 8B Instruct
Figure 9: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth (GT) dataset and Self-Output (SO) dataset of MATH task with Mistral 7B Instruct
Figure 10: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth (GT) dataset and Self-Output (SO) dataset of MATH task with Gemma 2 IT 2B
Appendix F Reproducibility
F.1 Training data cost
As we know there is efficiency difference of the data preparation between Self-Output, Rephrase, Selective Token Masking and Baseline Fine-tuning. To brief about the difference, Self-Output is the most resource-exhausted since it requires to generate $N$ samples (usually set 32), validate the correctness of samples (including parsing out the answer, and match with gold label or pass test cases), and then do forward passes on those samples to pick the lowest perplexity sentence as training data. On the other hand, STM only requires a single forward pass on ground truth data to calculate perplexity of each token to do the masking in training process. As Table 17 shows, for a simple MBPP task, on a A100 GPU server to prepare for 374 instances (train) + 90 instances (validation) for NN=32 takes about more than 18 gpu hours to complete the process of Self-Output to generate trainable data. While STM only requires 4 minutes to calculate the whole token perplexity.
@@ F.1 Training data cost
Baseline Fine-tuning 374 100% 0
STM Llama 3 8B Instruct 374 100% 4.5 minutes
Self-Output Llama 3 8B Instruct 213 57% 16.8 hours
Rephrase Llama 3 8B Instruct 327 87.4% 36.8 minutes
F.2 Model training resource
We train Llama 3 8B Instruct, Mistral 7B Instruct and Gemma 2 IT 2B using two NVIDIA A100 Tensor Core GPU with VRAM 40GB for each GPU and additional RAM of 96GB. In terms of gpu hour, Our training experiments takes at most 2 hours for MBPP task, 4 to 6 hours for MATH task. As for the evaluation experiments for other non-target tasks, evaluation in ARC-Challenge takes 3.5 hours, evaluation in GSM8k takes 4.5 hours, and evaluation in BIRD task takes 7.5 hours.
F.3 Evaluation prompt
@@
MBPP prompt user prompt Please refer the given test cases and generate a python function for my problem. Make sure the written code is wrapped in code block : ```python <yourcode><yourcode>``` >>>>>> Problem:{text}>>>>>> Test Cases:{test_list}
@@
ARC Challenge prompt user prompt Answer the following multiple choice question. The last line of your response should be of the following format: ’Answer: $\mathdollar LETTER^{\prime}$ (without quotes) where LETTER is one of ABCD. Question: {question} Choices: {choices}
In GSM8k task, we add a Question-Answering task instruction and an answer format instruction before the ’question’ as the user role prompt. To evaluate the output, we directly parse the format of ’####<numberonlyanswer>\#\#\#\#<number\,\,only\,\,answer>’ from the generated results, and match it with the ’answer’ value from the dataset. The TextBox shows an example of the user prompt:
GSM8k prompt user prompt You are given a grade school math question. Please answer the question in the following format: Q: <<question>> A: <<Think step by step here>> #### <<number only answer>> Format requirements : you must first output your reasoning before finalized with the " #### " format followed by the final numeric answer.
@@
Appendix G Other STM alternatives
@@
Original model 0% 28.1 51.1 58.5 74.7 13.9
Basline 0% 21.7 45.8 19.0 29.5 5.1
Fine-tuning
STMSTM 23.8% 30.3 49.5 55.4 75.3 13.3
@@
Appendix H A continual learning setting of STM
Table 19: STM applied on continual learning settings
steps Training Task MATH-test MBPP-test GSM8K ARC BIRD
0 Initial model 0.3140 0.5979 0.7255 0.7816 0.2053
1 MATH-train 0.334 0.6164 0.7687 0.7577 0.2027
2 MBPP-train 0.332 0.625 0.7635 0.7688 0.2073
Appendix I Does self-output response correctness matters?
In our experiment, the lowest perplexity scored data are collected exclusively from correct Self-Output subsets. In MBPP training task, a correct Self-Output data refers to a generated response that passes all the test cases. In MATH, correctness refers to a correct final answer parsed from a response. Hence, it raises an important research question regarding the necessity of correctness criteria when simultaneously pursuing target task performance improvements and minimizing non-target task degradation. In Table 20, we collect different correctness rates of Self-Output response as MBPP training datasets respectively with Llama 3 8B Instruct model. Correctness rate refers to the ratio of the samples whose labels are verified as "correct". Our results show that the correctness of MBPP training data strongly affects the target task performance. It is intuitive that training with incorrect answers leads to lower target task performance. However, compared to model trained on ground truth data (the correctness rate is 100%), which fails to sustain non-target task performance, models trained with higher correctness rate of Self-Output data achieves higher target task improvements and still sustains the non-target task performance robustness at the same time. Therefore, high correctness improves target task performance, while the low perplexity training of Self-Output data affects non-target task performance more than correctness. We have also conducted the correctness test with smaller models like Gemma 2 IT 2B Instruct, please refer to Appendix I.
Table 20: The performance of models trained with different correctness rate of MBPP Self-Output dataset and trainable tokens number. Values of a cell block in gray represents the target task testing performance of the Llama 3 8B Instruct trained with target task data.
@@ Appendix I Does self-output response correctness matters?
[…]
Self-output
[…]
We have also tested the ablation test of correctness with Gemma 2 IT 2B model as Table 21 shows. It holds the same conclusion that correctness is relevant to target task performance only, while the low perplexity trait of Self-Output data is more relevant to non-target task performance than correctness. However in Gemma 2 IT 2B setting, the number of training instance with at least one positive and one negative responses at the same time is much fewer than ground truth data and Llama 3 8B Instruct’s correctness data although it has more trainable tokens due to different tokenizers are applied. That is to say, to maintain the same training data amount between each correctness rate dataset for comparison in the Figure, we have to choose a smaller training data size for Gemma 2 IT 2B. Therefore, the effectiveness of data correctness is not as obvious as Llama 3 8B Instruct model does.
Table 21: The performances of models trained with different correctness rate of MBPP Self-Output dataset and trainable tokens number. Values of a cell block in gray represents the target task testing performance of the Gemma 2 IT 2B trained with target task data.
rejection rate Trained tokens MBPP MATH ARC GSM8k BIRD
[…]
Self-Output
[…]
Appendix J Regularization effects on non-target tasks
We apply a range for searching best combination of weight decay and dropout rate hyperparameters when training an MBPP task on Llama 3 8B Instruct model. The weight decay is set from 0.0 to 0.5 with interval 0.1, and dropout rate is set from 0.05 to 0.5 with an interval of 0.05 to 0.1. And we pick the top 3 best performing weight decay value with its dropout rate, and choose the best 2 performances among the checkpoints of 50 steps to 300 steps (about 5 to 6 epochs training) as our results. Our experiment results as Figure 11 show that, when applying dropout or weight decay on ground truth data, the model trained with target task like MBPP still overfits on training data and the testing performance degrades. On the other hand, the non-target performance of MATH as Figure 12 shows, we found that some of the results perform better than the original Llama 3 8B Instruct in non-target task, but their target task performances are found worse than the base model’s performance. It implies that the widely adopted regularization techniques are hard to optimized for both target task and non-target task performance.
Figure 11: The target task performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama 3 8B Instruct) and comparison between the original Llama 3 8B Instruct model’s performance on MBPP testing dataset.
Figure 12: The non-target MATH task performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama 3 8B Instruct) and comparison between the original Llama 3 8B Instruct model’s performance on MATH testing dataset.
Appendix K Rank analysis on LoRA weights
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We propose a novel empirical explanation for non-target task performance degradation, introducing Selective Token Masking (STM), a simple yet effective strategy that filters out high-perplexity tokens to enable low-perplexity training. We have applied STM on (1) different perplexity criteria, (2) different model families and scales, and (3) various fine-tuning techniques from full-weight fine-tuning to parameter-efficient fine-tuning. We found that STM could achieve almost the same capabilities as the original models have on non-target tasks after fine-tuning on target tasks under all circumstances. We could identify that the degradation results from the low model weights update during the low perplexity training. Together, these insights provide a new perspective on designing fine-tuning strategies that better retain general capabilities while enhancing performance on domain-specific tasks. Please refer to Section 5 for our results and the Appendix for more results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please refer to the last section after the conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We have our assumption made and found in Introduction and section 2 as a initiatives/motivations of our work. While it’s a more empirical research instead of theoretical research, so we do not cover the theory behind the models.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have mentioned where each baselines, dataset, metrics are adopted from. And explain how to implement our method as it’s very simple to understand and implement.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, we have provide our data and code within a github link in Section 4.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We mentioned our data curation in Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have added variance on loss error curve and perplexity calculations. All LLM metrics are done in a greedy (deterministic) way to avoid randomness noise involved.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We mentioned the training resource in Appendix F
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, the author has read through the entire Neurips Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We mentioned out work could bring positive impact on avoiding potential forgetting brought by fine-tuning. This helps the model training become more robust, and makes model safety and security issue more easily be handled.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work propose a training techniques to avoid catastrophic forgetting. Thus there’s no such risk of realeasing model or dataset that have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited all the public dataset, open-source models, closed-form API model that are used in our work.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: There is no new assets introduced in the paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This experiments conducted in this paper did not involve any human subjects experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper is mainly a study in Fine-tuning on Large Language Models which has no human or any biological participants involved.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The use of LLMs is only limited in Latex syntax queries.
  80. Guidelines: This work falls under the field of LLMs (Neural Network) finetuning and do not involve any experiments which contain living participants.
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.